OpenStudy (richyw):

Consider a particle in a potential \(V(x)\) represented by a definite energy wave function, sollution of the corresponding time-independent Shrodinger equation given by \[\Psi_n(x,t)=\psi_n(x)e^{-i\omega_nt}\]where \(E_n=\hbar \omega_n\) Using the energy operator \[\hat{E}_n=i\hbar \frac{\partial}{\partial t}\]

OpenStudy (richyw):

a) obtain the expectation value of E, E;

OpenStudy (richyw):

This is an assignment question so please don't just give me the answer. But some help would be very much appreciated!

OpenStudy (anonymous):

hate quantum . . .

OpenStudy (richyw):

I am just getting started on it (obviously!)

OpenStudy (fwizbang):

Is the wavefunction in an eigenstate of energy?

OpenStudy (richyw):

to be honest. I don't even know what that means. I found out the answer once I knew that\[\int^\infty_{-\infty}\Psi^*(x)\Psi(x)=1\]

OpenStudy (fwizbang):

Certainly one way to get the answer is by direct calculation. the expectation value of any operator is \[< O(t) > =\int\limits_{-\infty}^{\infty}\psi^*(x,t)O \psi(x,t) dx\] where O is the operator. Another method(in this case at least) is to notice that the wvfct is an energy eiegenstate, which means that \[\hat E \Psi = \hbar \omega_n \Psi\] which means that for this wvfct, the only possible result of a measurement of E would be \[ \hbar\omega_n \].